Fermi National Laboratory
 

Accelerator Update for June 30, 2006 - July 5, 2006

 

Friday June 30

The day shift began with Operations monitoring store 4804 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam store, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, Main Injector (MI), and Recycler operating normally, and with MiniBooNE and NuMI taking beam.

Operations terminated store 4804 at 8:03 AM.

Operations put the TeV into shot setup around 9:40 AM.

At 10:33 AM, the NuMI Horn tripped off, which also tripped off the NuMI beam permit. Operations contacted experts.

Operations established store 4805 at 10:55 AM with an initial luminosity of 142.5E30. Pbar resumed stacking.

Operations began sending beam to the Meson MTest experiment at 1:42 PM.

The recycler began preparing stash antiprotons at 3:27 PM. The stash ended at 4:24 PM. Pbar and MTest resumed taking beam.

At 5:18 PM, NuMI experts reported a problem with the horn LCW system flow. Experts continued to investigate and at 8:38 PM reported that a deionizing bottle’s resin beads broke apart and clogged up the horn. Experts reported that the experiment would be off for several days.

The Recycler began stashing at 9:17 PM. The stash ended at 10:51 PM.

Saturday July 1

The midnight shift began with Operations monitoring store 4805 and the stash, with Pbar stacking, and with MTest and MiniBooNE taking beam.

The MTest experiment shut down for the night at 12:03 AM.

The Recycler began stashing at 5:46 AM. The stash ended at 7:02 AM.

At 5:50 AM, the Booster West Anode Power Supply (WAPS) tripped off for the first of many times over the next 75 minutes. Operators contacted an expert who suggested a timeline change.

At 8:39 AM, store 4805 aborted. There was no quench.

At 8:40 AM, a Booster quadrupole power supply tripped off and wouldn’t reset. An operator called in an expert and then later another expert came in to assist. They had the supply back up by 12:17 PM. Beam from Booster resumed at 12:26 PM.

The Recycler began stashing at 1:20 PM. The stash ended. Normal operations resumed around 2:21 PM.

Operations put the TeV into shot setup at 3:18 PM.

Operations established store 4808 at 6:01 PM, with an initial luminosity of 120.62E30. Normal beam operations resumed.

Sunday July 2

The midnight shift began with Operations monitoring store 4808 and the stash, with Pbar stacking, and with MTest and MiniBooNE taking beam.

The MTest experiment shut down for the night at 12:22 AM.

The Recycler began stashing at 6:11 AM. The stash ended. Stacking resumed at 6:41 AM.

At 4:41 PM, a Booster kicker tripped off. Operators called in an expert. He repaired a kicker cable. Booster beam resumed at 6:19 PM.

Operations terminated store 4808 at 10:03 PM.

Operations put the TeV into shot setup at 10:39 PM.

At 11:44 PM, the TeV suffered at sector E4 quench.

Monday July 3

The midnight shift began with Operations monitoring a stack and stash, with Operations investigating the quench, and with MiniBooNE taking beam.

Operations rebooted the TeV and MI low-level RF systems at 3:21 AM.

Pbar resumed stacking at 3:32 AM.

Operations rebooted the MI dampers at 3:42 AM.

Pbar resumed stacking at 5 AM.

At 5:06 AM, a Recycler power supply (R:CE304) tripped off. The Recycler lost some of its stash. Operations contacted an expert.

At 5:08 AM, Operations reported a problem with the TeV bus and sector B3. An expert came in and replaced a DC breaker. The TeV ramped up at 7:47 AM.

Operations put the TeV into a dry squeeze at 8:05 AM.

At 8:07 AM, Operations reported that a lightning storm caused several systems to trip.

MTest resumed taking beam at 9:15 AM.

Operations prepared CDF and D0 for access at 11:16 AM.

Operations put the TeV into shot setup at 3:10 PM.

Operations established store 4812 at 4:40 PM with an initial luminosity of 116E30. Normal operations resumed.

Tuesday July 4

The midnight shift began with Operations monitoring store 4812 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

The Recycler stashed at 3:37 AM and stacking resumed at 4:09 AM.

At 7:57 AM, the MI power supply loop tripped off on a ground fault indication.

Operations reported at 8:09 AM that they had completed a successful hipot of MI.

At 10:10 AM, Operations reported that a MI-40 power supply had a shorted SCR. Power supply experts replaced 5 SCRs and repaired a LCW leak.

Pbar resumed stacking at 1 PM.

The Recycler stashed and stacking resumed at 4:10 PM.

MTest quit taking beam at 7 PM.

Wednesday July 5

The midnight shift began with Operations monitoring store 4812 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

Operation terminated store 4812 2:02 AM and put the TeV back into shot setup fifteen minutes later.

Operations established store 4813 at 4:28 AM with an initial luminosity of 133.27E30. Normal operations resumed.

The Recycler stashed and stacking resumed at 6:26 AM.

The Plans for Wednesday

The plans for today are to monitor the store. Booster will be allowed a five-hour access to replace some RF power amps. Other machines, except for the Tevatron, will use this time to conduct maintenance work.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/05/2006    email Fermilab